A basketball player does 2.45×105 J of work during her time in the game, and evaporates 0.110 kg of water. Assume a latent heat of 2.26×106J/kg for the perspiration (the same as for water). Determine the number of nutritional calories the player has converted to work and heat.
Answer #1

Heat produced = mL where m is mass, L is latent heat for perspiration

= 0.11*2.26e6 = 248600 J

Number of nutritional calorie converted to work = 2.45e5/4186 Cal = 58.5 Cal

Number of nutritional calorie converted to heat = 248600/4186 Cal = 59.4 Cal